Prove the following trig identities

sin 2x + cos 2x / sin x + cos x = 2cosx

2sinxcosx / cos4x-sin4x = tan2x


1
Expert's answer
2021-09-10T00:14:19-0400

(b)


"\frac{2sinxcosx}{cos^4x-sin^4x}=tan2x"

taking LHS

"=\frac{sin2x}{(cos^2x+sin^2x)(cos^2x-sin^2x)}\\=\frac{sin2x}{cos^2x-sin^2x}\\=\frac{sin2x}{cos2x}\\=tan2x"

(a)

"\frac{sin2x+cos2x}{cosx+sinx}=2cosx"

taking LHS

"=\frac{2sinxcosx+2cos^2x-1}{cosx+sinx}"

"2cosx- \frac{1}{sinx+cosx}"

so sinx +cosx cannnot be zero so this is not solved further